OpenStudy (frank_11):

Why can -4 be an answer for |x-24|=7x

OpenStudy (trojanpoem):

I don't know either ... L.H.S = | -4 - 24| = 28 R.H.S = 7 * -4 = -28 L.H.S != R.H.S I need help too ...

TheSmartOne (thesmartone):

An absolute value cannot be equal to a negative value. |-4 - 24 |= |-28| = 28 -4 * 7 = -28 28 \(\neq\) -28

OpenStudy (agent0smith):

"Why can -4 be an answer for |x-24|=7x" My guess is there's a typo there...

OpenStudy (kevin):

|x - 24| = 7x We have two possibility answers : (x - 24) = 7x OR -(x - 24) = 7x For (x - 24) = 7x : x - 24 = 7x -24 = 7x - x -24 = 6x x = -24/6 x = -4 --------> This is the answer of your question But remember every time you done finding variables in an equation, you should check it on original equation |x - 24| = 7x |-4 - 24| = 7(-4) |-28| = -28 28 = -28 --> This is not correct So -4 is not the answer for |x-24|=7x

OpenStudy (kevin):

I think your question must be: Why can't -4 be an answer for |x-24|=7x ?

OpenStudy (mathmale):

I agree with Kevin's response (above), and that of TheSmartOne. The output of the absolute value function can NOT be negative; it must be ZERO or POSITIVE.